An electron that has a velocity with x component 2.0 x 106 m/s and y component 2.9 x 106 m/s moves through a uniform magnetic field with x component 0.032 T and y component -0.19 T. (a) Find the magnitude of the magnetic force on the electron. (b) Repeat your calculation for a proton having the same velocity. answer must be in N units!!!!!!
Answer #1

given

vx = 2*10^6 m/s

vy = 2.9*10^6 m/s

Bx = 0.032 T

By = -0.19 T


a) magnetic force, F = q*(v cross B)

= q*( (2*10^6i + 2.9*10^6j) cross (0.032 i - 0.19 j))

= q*(2*10^6*(-0.19) k + 2.9*10^6*0.032*(-k))

= q*(-0.4728*10^6) N

= (-1.6*10^-19)*(-0.4728*10^6) N

= 7.56*10^-14 N (towards +z axis)

b) same. i.e 7.56*10^-14 N (towards -z axis)